# The Magnetic Filling in MAD Simulations and its Impact on the Jet in M87

### v1

@@ Appendix D MRI Quality Factors
Figure 18: Disk averaged MRI quality factors in $\theta$ and $\phi$ of each simulation, rowwise. Dashed lines are the median values taken over the time span shown in the graphs. Besides the median values we also give corresponding time averaged values. Due to the scewed nature of the temporal distriubtion of the disk averaged Q-factors, the median is more appropiate for interpretation. It is noted that for $Q^{(\theta)}$ $\sim(3-6)\%$ of points lie outside the graph, i.e. are above $100$ in the three simulations. For $Q^{(\phi)}\sim 1\%$ are concerned.
Figure 19: Meridional cross-section of MRI Quality factors in $\theta$ and $\phi$ directions. Top row: Meridional cross-section. Bottom row: Volume averaged Q-factors over the full azimuthal angle. The black lines indicate the regions where the disk is cut, poloidally, and subsequently averaged over the full azimuthal angle to obtain the disk averaged values given in the bottom left corners.
@@ Appendix E Error Estimation of Time Series
The autocorrelation function is computed up to a lag of $N/2$ to avoid noise-dominated estimates at large lags and normalized to lie between values of 1-1 and $1$ to anable comparability between the simulations.
@@
where $\sigma$ is the standard deviation of the time series. We typically find values of $\tau_{\rm{auto}}\lesssim 10M$ corresponding to an error $\lesssim 5$ higher than the standard error of the mean, when assuming the data is entirely uncorrelated. Therefore, we avoid underestimating the error of the mean.

### v2

@@ Appendix D MRI quality factors
Figure 18: Disk averaged MRI quality factors in $\theta$ and $\phi$ of each simulation, row-wise. Dashed lines are the median values taken over the time span shown in the graphs. Besides the median values we also give corresponding time averaged values. Due to the skewed nature of the temporal distriubtion of the disk averaged $Q$ factors, the median is more appropriate for interpretation. It is noted that for $Q^{(\theta)}$ $\sim(3-6)\%$ of points lie outside the graph, i.e. are above $100$ in the three simulations. For $Q^{(\phi)}\sim 1\%$ are concerned.
Figure 19: Meridional cross-section of MRI Quality factors in $\theta$ and $\phi$ directions. Top row: Meridional cross-section. Bottom row: Volume averaged $Q$ factors over the full azimuthal angle. The black lines indicate the regions where the disk is cut, poloidally, and subsequently averaged over the full azimuthal angle to obtain the disk averaged values given in the bottom left corners.
@@ Appendix E Error estimation of time series
The autocorrelation function is computed up to a lag of $N/2$ to avoid noise-dominated estimates at large lags and normalized to lie between values of 1-1 and $1$ to enable comparability between the simulations.
@@
where $\sigma$ is the standard deviation of the time series. We typically found values of $\tau_{\rm{auto}}\lesssim 10M$ corresponding to an error $\lesssim 5$ higher than the standard error of the mean, when assuming the data are entirely uncorrelated. Therefore, we avoid underestimating the uncertainty of the mean.